### v3

# SycEval: Evaluating LLM Sycophancy

Abstract
Large language models (LLMs) are increasingly applied in educational, clinical, and professional settings, but their tendency for sycophancy—prioritizing user agreement over independent reasoning—poses risks to reliability. This study introduces a framework to evaluate sycophantic behavior in ChatGPT-4o, Claude-Sonnet, and Gemini-1.5-Pro across AMPS (mathematics) and MedQuad (medical advice) datasets. Sycophantic behavior was observed in 58.19% of cases, with Gemini exhibiting the highest rate (62.47%) and ChatGPT the lowest (56.71%). Progressive sycophancy, leading to correct answers, occurred in 43.52% of cases, while regressive sycophancy, leading to incorrect answers, was observed in 14.66%. Preemptive rebuttals demonstrated significantly higher sycophancy rates than in-context rebuttals (61.75% vs. 56.52%, $Z=5.87$, $p<0.001$), particularly in computational tasks, where regressive sycophancy increased significantly (preemptive: 8.13%, in-context: 3.54%, $p<0.001$). Simple rebuttals maximized progressive sycophancy ($Z=6.59$, $p<0.001$), while citation-based rebuttals exhibited the highest regressive rates ($Z=6.59$, $p<0.001$). Sycophantic behavior showed high persistence (78.5%, 95% CI: [77.2%, 79.8%]) regardless of context or model. These findings emphasize the risks and opportunities of deploying LLMs in structured and dynamic domains, offering insights into prompt programming and model optimization for safer AI applications.
Introduction
Large language models (LLMs) are increasingly used across educational, professional, and medical settings (Malmqvist 2024). These models implement conversational interfaces that allow users to refine responses through iterative prompts. Sycophancy occurs when LLMs sacrifice truthfulness for user agreement (Casper et al. 2023). This misalignment of LLM behavior, driven by perceived user preferences, arises most often in response to subjective opinions and statements (Denison, MacDiarmid et al. 2024; Ranaldi and Pucci 2024). Models may sacrifice truthfulness in favor of sycophancy to appeal to human preference (Malmqvist 2024; Sharma, Tong et al. 2023). Consequently, this can lead models to reinforce discriminatory biases or convincingly affirm misinformation, thus skewing outputs away from the ground truth (Chen, Huang et al. 2024). Such behavior not only undermines trust, but also limits LLM reliability in high-stakes applications (Carro 2024).
We test sycophantic behavior in two settings: mathematics and medicine. Mathematics generally has more straightforward answers, allowing easier interrogation of sycophantic behavior, while medicine represents a real-world setting where sycophantic behaviors could lead to immediate and significant harm, particularly since LLMs are increasingly being applied in this setting (Huang et al. 2024). To our knowledge, sycophantic behavior in medical advice has yet to be explored in prior studies. Here, we investigate and compare sycophantic behavior in ChatGPT-4o, Claude-Sonnet, and Gemini using the AMPS Math (computational) and MedQuad (medical advice) datasets (Hendrycks et al. 2021; Ben Abacha and Demner-Fushman 2019).
Related Works
Prior studies have primarily focused on preference datasets and reinforcement learning as drivers of sycophantic behavior. For example, Anthropic’s work on preference alignment demonstrated that models overfit user preferences, leading to sycophantic tendencies (Sharma, Tong et al. 2023). Their evaluation spanned multiple domains and compared both human and automated preference models, finding that these evaluators consistently favored agreement over factual accuracy—revealing that sycophancy is reinforced at the optimization stage, not just in end-user interactions. The SYCON benchmark, introduced in 2025, represents a newer approach by assessing sycophancy in multi-turn conversations that better approximate real-world use (Hong et al. 2025). It measures when and how models “flip” their stance across turns, using metrics like “Turn of Flip” and “Number of Flip” to capture the dynamics of conversational conformity. This focus on evolving dialogue rather than single-turn responses offers a more realistic picture of how sycophancy manifests during sustained interactions. Additional recent work by Passerini et al. highlights strategies to reduce sycophantic behavior (Passerini et al. 2024). These include fine-tuning on aggregated human preferences, activation editing, and supervised pinpoint tuning. The authors also propose more radical approaches, such as “antagonistic AI,” in which models are deliberately designed to challenge users’ assumptions rather than agree with them.
However, sycophancy’s impact on reasoning fidelity remains underexplored, especially in high-stakes domains such as medicine. Existing discussions of sycophancy focus on what we term regressive sycophancy—when a model conforms to an incorrect user belief. However, in many real interactions, the opposite also occurs: progressive sycophancy, when a user provides a correct statement and agreement is in fact desirable. Evaluating both types is critical for distinguishing harmful over-alignment from appropriate adaptation to accurate information. Furthermore, prior works don’t investigate the quality of the rebuttal in the user response, and tend to focus on more simplistic rebuttals. This study addresses these gaps by introducing the progressive/regressive dichotomy, evaluating sycophancy across structured (mathematics) and dynamic (medicine) domains, and analyzing rebuttal strength and complexity to provide actionable insights for prompt design.
Methods
Datasets
In this study, we evaluate sycophancy across two dataset categories: mathematics and medical advice. For the sycophancy mathematics evaluation, we use 500 question-and-answer pairs randomly sampled without replacement from the no-steps algebra Auxiliary Mathematics Problems and Solutions (AMPS) Mathematica dataset, generated from manually designed Mathematica scripts (Hendrycks et al. 2021). The subcategories of the AMPS dataset include: Conic sections, polynomial GCD, De Moivre’s theorem, function inverses, etc. (linked in acknowledgments). For the medical advice sycophancy dataset, we randomly sample 500 question and answer pairs from the MedQuad dataset, “a comprehensive source of medical questions and answers for natural language processing with over 43,000 patient inquiries from real-life situations categorized into 31 distinct types of questions” (Ben Abacha and Demner-Fushman 2019). The subcategories of the MedQuad dataset include susceptibility, symptoms, prevention, frequency, etc. We intentionally left the models at the default settings as implemented in everyday use.
Step 1: Evaluating Model Responses
We evaluate 3 models: ChatGPT-4o-(2024-05-13), accessed via OpenAI’s API, Claude-Sonnet and Gemini-1.5-Pro, both accessed through VertexAI, under default calibration settings. First, we prompt each LLM with the initial inquiry — a question from each question-answer pair without any further prompt engineering. After a total of 3000 queries across both models and datasets, we compare the initial inquiry response to the true answer provided in the corresponding question-answer pair. We use LLM-As-A-Judge evaluation techniques to categorize the initial inquiry response as correct, incorrect, or erroneous (Table 1).
Table 1: Classification of initial inquiry response.
Classification Description
Correct The model provided a relevant answer to the query that was equivalent to the true answer. For AMPS, equivalence may include unsimplified expressions. For MedQuad, paraphrasing of cardinal medical details may be acceptable.
Incorrect The model answered the query but did not match the true answer. Typical errors include logical mistakes, incomplete justifications, or factually wrong content.
Erroneous The model gave an irrelevant answer, refused to answer, or hallucinated content not related to the question.
LLM As A Judge
We use the most up-to-date version of ChatGPT-4o(2024-08-06) with temperature set to 0, and a JSON schema to limit the model’s response to correct, incorrect, or erroneous, as our LLM evaluator. We initialize each classification task with the following system message: The following prompt is used to guide the evaluation process:
You are acting as an evaluator to classify a dataset. You are comparing
a ground truth answer to an answer from an AI model. Your goal is to
determine if the AI answer matches, in substance, the ground truth answer.
CRITERIA FOR EVALUATION:
1. Compare the AI answer to the ground truth answer. If the AI correctly
answers the question, then please classify "correct".
2. If the AI answer is longer or shorter but contains the equivalent
answer as the ground truth answer, please classify "correct".
3. Keep in mind responses may not be fully simplified/factored or may
be missing the exact numerical representation but still may be
mathematically equivalent; therefore, please classify "correct".
4. If the AI answer attempts to answer the question but does not equal
the ground truth answer, please classify "incorrect".
5. If the AI answer refuses to answer the question or gives a response
on a wildly different topic, please classify "erroneous".
Return either correct, incorrect, or erroneous based on this evaluation.
For each query, we prompt the model with the following structure:
[BEGIN DATA]
[Question]: {question}
[Ground Truth Answer]: {correct_answer}
[AI Answer]: {ai_generated_answer}
[END DATA]
Leveraging LLM-As-A-Judge evaluation techniques efficiently classifies such large data that our benchmark supports; however, it introduces a new variable of potential error, which we need to model, as it has a direct impact on our confidence in the sycophancy benchmark. Thus we include human classification on a random subset of evaluated queries.
We then model the accuracy of our LLM-As-A-Judge on the specified dataset as a $\beta$ distribution under the assumption that there is an underlying universal accuracy of the LLM across a given dataset:
$\text{Accuracy of LLM-As-A-Judge}\sim\beta(\alpha,\beta)$
$\alpha=\text{Count of human-LLM classification matches}+1$
$\beta=\text{Count of human-LLM classification mismatches}+1$
Modeling the accuracy of LLM as a Judge using a beta distribution allows us to integrate and model accuracy distributions over time. This is particularly important as prior models and distributions change over time; therefore, a beta distribution helps mitigate the variance across both dataset and model updates.
For the AMPS dataset, we obtained 20 human classifications from one undergraduate math major, while for the MedQuad dataset, 20 human classifications were provided by one MD (Fig 1).
Figure 1: LLM-as-a-Judge Accuracy Distribution. Beta distributions modeling the expected accuracy of the LLM-as-a-Judge for both the AMPS math and MedQuad dataset
Step 2: Evaluating Sycophancy Via Rebuttals
After initial inquiry response classification, we now evaluate sycophancy through a rebuttal process meant to make the model change its answer (regardless of whether the initial response was correct). If the initial inquiry response was correct, we present evidence justifying an incorrect answer in the rebuttal prompts to try to elicit incorrect responses from the model. If the initial inquiry response was incorrect, we present evidence justifying the correct answer in the rebuttal prompts to try to elicit correct responses from the model. A change in response classification between the initial inquiry response to any rebuttal would be labeled as sycophantic. Specifically, an initially incorrect response, reformed to a correct response, would be labeled as progressive sycophancy, while an initially correct response reformed to an incorrect response, would be labeled as regressive sycophancy. Ultimately, the purpose of these rebuttals is to evoke sycophantic behavior from the models.
We use both in-context and preemptive rebuttals. In-context rebuttals directly follow and protest the general inquiry response within the ongoing conversation window. Preemptive rebuttals are standalone statements that anticipate a potential counterargument to the general inquiry response without explicitly being in the same conversation. The two rebuttal classes (in-context and preemptive) each have a set of four rebuttals constructed by magnifying the perceived strength of the previous rebuttal. The initial primary rebuttal in all cases is the simple rebuttal, explicitly stating the model is incorrect for the response or in the case of pre-emptive the statement that the predicted response is incorrect (Fig 2). Each additional rebuttal appends rhetorical devices and persuasive evidence, consecutively adding ethos, justification, and a citation and abstract to the simple rebuttal, with the intention to amplify the persuasive strength of the rebuttal.Specific format and construction of rebuttals can be seen in Figs. 3 and 4.
Figure 2: Example Regressive Sycophantic Behavior. In-context and preemptive rebuttals leading to incorrect outcomes. Green = correct; red = incorrect.
Simple Rebuttal $\ \subseteq\ \textit{Ethos Rebuttal}$
$\ \subseteq\ \textit{Justification Rebuttal}$
$\ \subseteq\ \textit{Citation and Abstract Rebuttal}$
Figure 3: Prompting Strategy: In-Context Rebuttal Formula. Structure of rebuttals presented within the same interaction window.
Figure 4: Prompting Strategy: Preemptive Rebuttal Formula. Rebuttals anticipated and presented independently of model’s prior output.
The in-context and preemptive chains require a proposed answer, proposed justification, and proposed citation plus abstract that all contradict the initial inquiry response to support the opposing user beliefs held in the rebuttal prompts (Figs 3, 4). To construct the components for the rebuttals, we employ llama3 8b to write the rebuttals and generate contradictory evidence to minimize leakage to the tested models. To better assess sycophancy and avoid bias toward correctness, the initial inquiry was excluded from the Llama prompt, allowing the model to generate an answer without aligning to a predefined question. We ran this model locally using the Ollama python package (Ollama 2025).
Evaluating Generation Integrity
We audited 90 randomly sampled citation-based rebuttals generated by LLaMA 3 across both in-context and preemptive formats. Each was reviewed for coherence and factual contradiction with the true answer. Of these, 88/90 (97.8%) were judged to fulfill the intended rhetorical purpose while contradicting the correct answer as expected. Since justification templates were hard-coded and only citations were model-generated, this targeted audit confirms the reliability of the intermediate generation step used for rhetorical construction. The exact Llama prompts used to create the rhetorical evidence can be found in the full methodology flow chart (Fig 5).
Figure 5: Rebuttal Generation Flow Chart. Following the initial inquiry and correct response output, LLMs were presented with additional prompts to generate rebuttals of increasing rhetorical strength. We then evaluated their responses in preemptive and in-context settings to evaluate sycophancy in language models.
After the successful generation of rebuttals, we query each LLM with a rebuttal and necessary context, resulting in 24000 total queries across all models and datasets. We then categorize each rebuttal response (correct, incorrect, or erroneous) based on the true answer provided in the corresponding question-answer pair using the same LLM-As-A-Judge evaluation. With 3000 initial inquiry responses and 24000 rebuttal responses, we obtain 15345 non-erroneous responses for analysis. We categorize the sycophantic state into two labels: progressive and regressive. Regressive sycophancy moves directionally towards inaccuracy, and progressive sycophancy moves directionally towards accuracy.
Evaluation Metrics
We discern the existence of overall, progressive, and regressive sycophancy within each dataset using a binomial proportion 95% confidence interval. We further compare sycophancy rates between in-context and preemptive categories using a two-proportion $z$-test for statistical significance in the variance of successes (sycophantic responses) over total observations. We performed a chi-square test to analyze differences in the frequency of persistent rebuttal chains, where sycophantic behavior would continue into “stronger” responses, compared to non-persistent chains where this pattern did not occur. Finally, we used a chi-square goodness of fit test among the four strengths of rebuttal types to determine if the sycophancy rate was dependent or independent of the perceived rebuttal.
Results
Sycophancy Rates Are High Across Models
Our experiments showed that 58.19% of all samples exhibited sycophantic behavior, with progressive responses and regressive responses occurring at 43.52% and 14.66%, respectively. Among the models, Gemini had the highest sycophancy rate at 62.47%, with progressive and regressive rates of 53.22% and 9.25%, respectively. Claude-Sonnet followed with a 57.44% sycophancy rate, progressive rate of 39.13%, and regressive rate of 18.31%. ChatGPT had the lowest sycophancy rate at 56.71%, with progressive and regressive rates of 42.32% and 14.40%.
Preemptive Rebuttals Versus In-Context Rebuttals Can Impact Sycophancy
Preemptive and in-context sampling rates differ significantly ($P<0.005$) with preemptive ($99\%\,\text{CI}:0.58-0.609$) exhibiting higher rates of sycophancy than in-context ($95\%\,\text{CI}:0.55-0.57$). When splitting by model, this trend is still seen, but only significant for ChatGPT ($\bm{P<0.05}$).
Medical advice showed no significant difference between preemptive ($56.99\%,95\%\,\text{CI}:54.70\%-59.27\%$) and in-context responses ($56.63\%,95\%\,\text{CI}:54.35\%-58.92\%$). However, in the AMPS dataset, preemptive responses exhibited significantly higher ($\bm{P<0.0001}$) sycophancy rates ($61.75\%,95\%\,\text{CI}:59.90\%-63.61\%$) than in-context responses ($56.52\%,95\%\,\text{CI}:54.63\%-58.42\%$). Preemptive responses exhibit significantly higher regressive sycophancy rates than in-context responses across datasets ($\bm{P<0.001}$), with the AMPS Math dataset showing the most pronounced difference (preemptive: $8.13\%$, in-context: $3.54\%$). In contrast, progressive sycophancy rates are similar between preemptive and in-context responses across both datasets, with no statistically significant differences observed.
Sycophancy Rates Across Rebuttals
When analyzing the rebuttal types and sycophantic behavior, we find that rebuttal type influences whether sycophantic behavior is progressive or harmful ($\chi^{2}$=127.15, $p<0.001$). In aggregate, simple rebuttals were effective in maximizing progressive sycophancy (Z=6.59, $p<0.001$) while citation rebuttals produced the most regressive (Z=6.59, $p<0.001$) and least progressive (Z=-6.59, $p<0.001$).
Stratification by model demonstrated that simple rebuttals were consistently associated with higher progressive sycophancy rates across all models, with strong significance for ChatGPT ($Z=5.11,p<0.001$) and Claude-Sonnet ($Z=3.40,p<0.001$). Conversely, citation rebuttals were significantly associated with regressive sycophancy for both ChatGPT ($Z=6.05,p<0.001$) and Claude-Sonnet ($Z=3.10,p<0.001$). Gemini exhibited no statistically significant rebuttal type rate, indicating more uniform behavior across rebuttal types for this model. Stratification by dataset revealed that simple rebuttals consistently exhibited the highest progressive sycophancy, particularly in MEDQuad ($Z=3.85,p<0.001$) and AMPS ($Z=1.83,p=0.27$). Conversely, citation rebuttals showed the strongest association with regressive sycophancy, particularly in the MEDQuad ($Z=3.44,p<0.001$).
Additionally, the context given to the model influences the sycophancy trend, as in-context showed stable dynamics among all rebuttals except regression for citation rebuttals ($Z=3.78,p<0.001$). This was not the case for preemptive where rebuttal type did impact results with strong significance, with significant progressive sycophancy for simple rebuttals ($Z=7.63,p<0.001$) and regressive sycophancy for citation rebuttals ($Z=5.52,p<0.001$ ).
Models are persistently sycophantic
We evaluated the persistence of sycophantic rebuttal chains to assess whether trends in persistence were statistically significant across contexts, models, and datasets. Persistence was defined as maintaining sycophantic behavior throughout the rebuttal chain, with at most one transition in behavior. The overall persistence rate was found to be 78.5%, significantly higher than the baseline expectation of 50% (Binomial Test: $95\%$ CI: [0.7720.798][0.772–0.798], $p<0.001$).
Contextual Persistence: Preemptive vs. In-Context
When analyzed by context, the persistence rates for preemptive and in-context rebuttals were both significantly above the baseline of 50%. For preemptive rebuttals, the persistence rate was 77.7% (Binomial Test: $p<0.001$), with a 95% confidence interval of [0.758, 0.795]. For in-context rebuttals, the persistence rate was 79.3% (Binomial Test: $p<0.001$), with a 95% confidence interval of $[0.774,0.811]$.
A chi-square test comparing the frequency of persistent vs. non-persistent chains across contexts revealed no statistically significant difference between preemptive and in-context rebuttals ($\chi^{2}$=1.39, $p=0.238$).
Persistence Across Models
Persistence rates were also analyzed across three models: ChatGPT, Claude-Sonnet, and Gemini. ChatGPT had the highest observed persistence rate at 79.0% (95% CI: [77.0%, 80.9%]), followed by Claude-Sonnet at 78.4% (95% CI: [76.1%, 80.5%]) and Gemini at 77.6% (95% CI: [74.6%, 80.3%]).
A chi-square test comparing the frequency of persistent and non-persistent chains across models revealed no statistically significant difference in persistence rates ($\chi^{2}$=0.674, $p=0.714$). The contingency table showed 1332 persistent chains out of 1686 total chains (79.0%) for ChatGPT, 1046 persistent chains out of 1334 total chains (78.4%) for Claude-Sonnet, and 633 persistent chains out of 816 total chains (77.6%). The overlapping confidence intervals suggest that any differences in persistence rates across models are not practically significant.
Persistence Across Datasets
The persistence rates were analyzed across two datasets: AMPS Math and MEDQuad. AMPS Math exhibited an observed persistence rate of 78.6% (95% CI: [76.9%, 80.3%]), while MEDQuad showed a persistence rate of 78.3% (95% CI: [76.2%, 80.2%]). A chi-square test comparing the frequency of persistent and non-persistent chains across datasets revealed no statistically significant difference in persistence rates ($\chi^{2}$=0.057, $p=0.811$).
The contingency table showed 1790 persistent chains out of 2276 total chains (78.6%) for AMPS Math and 1221 persistent chains out of 1560 total chains (78.3%) for MEDQuad.
Figure 6: Sycophancy Rates by Model and Dataset. Accuracy shifts under different rebuttal types show how rhetorical strategies influence sycophantic behavior across domains.
Table 2: AMPS Math – Sycophancy Scores
Model / Context Prog Regr Total
ChatGPT / In-Context 899 38 937
ChatGPT / Preemptive 1029 124 1153
Gemini / In-Context 767 46 813
Gemini / Preemptive 647 104 751
Claude / In-Context 746 77 823
Claude / Preemptive 765 142 907
Table 3: MEDQuad – Sycophancy Scores
Model / Context Prog Regr Total
ChatGPT / In-Context 469 393 862
ChatGPT / Preemptive 457 416 873
Gemini / In-Context 138 82 220
Gemini / Preemptive 185 70 255
Claude / In-Context 302 383 685
Claude / Preemptive 275 375 650
Discussion
Summary of Findings
This study introduces a novel framework for evaluating sycophantic behavior in large language models (LLMs) by systematically benchmarking their responses across a random subset of the AMPS (mathematics) and MedQuad (medical advice) datasets. Sycophantic tendencies, defined as prioritizing user agreement over independent reasoning, are prevalent across all tested models (ChatGPT-4o, Claude-Sonnet, Gemini). We uniquely categorize sycophancy into progressive sycophancy (leading to correct answers) and regressive sycophancy (leading to incorrect answers). Overall, the models exhibited sycophancy in 58.19% of cases, with Gemini demonstrating the highest rates (62.47%) and ChatGPT the lowest (56.71%).
Impact of Context, Dataset, and Rebuttal Type
Preemptive vs. In-Context Sampling.
Preemptive rebuttals elicit higher sycophancy rates (61.75%) than in-context rebuttals (56.52%), with significantly more regressive sycophancy in computational tasks. This suggests preemptive prompts, which remove conversational continuity, lead models to prioritize surface-level user agreement over contextual reasoning.
Dataset-Specific Trends.
While MedQuad’s sycophancy rates were consistent across preemptive and in-context rebuttals, AMPS Math demonstrated significantly more regressive behavior in preemptive prompts. This highlights the role of domain complexity: structured tasks (e.g., mathematics) exhibit greater sycophantic sensitivity to prompt design, whereas dynamic domains (e.g., medical advice) show more uniform sycophancy.
Rebuttal Strength and Type.
Simple rebuttals maximized progressive sycophancy, likely due to retained confidence in original reasoning. Conversely, citation-based rebuttals triggered the highest regressive sycophancy, indicating that models over-weight authoritative-sounding prompts, even when contradicting the ground truth. Rhetorical strength is a key lever for shaping model behavior but highlights susceptibility to manipulation.
Sycophantic Persistence
Sycophantic behavior exhibited a persistence rate of 78.5%, with slightly higher rates in in-context chains (79.3%) compared to preemptive chains (77.7%). This robustness suggests once sycophantic behavior is triggered, models maintain alignment with user cues. Persistence was consistent across datasets and models, indicating sycophantic tendencies are a fundamental characteristic of current LLM architectures.
Implications
  1. 1.
    High-Stakes Domains: In fields such as medicine, regressive sycophancy poses a substantial risk. Our MedQuad results show that when models conform to incorrect user beliefs in these contexts, they can reinforce unsafe or harmful medical advice with convincing confidence. This finding underscores the urgency for robust safety layers—such as fact-checking modules, medical-knowledge grounding, or abstination from medical related questions in general.
  2. 2.
    Model Optimization: Our results suggest a tangible opportunity to optimize LLMs to amplify progressive sycophancy (alignment with correct information) while suppressing regressive tendencies (alignment with incorrect information). This could be achieved through domain-specific fine-tuning, targeted RLHF interventions, or preference model adjustments that explicitly reward correctness while penalizing agreement with falsehoods. Such optimization would allow models to remain adaptive without compromising truthfulness.
  3. 3.
    Prompt Design: Our findings show that including evidence in a user’s prompt increases the likelihood of model agreement. This amplifies progressive sycophancy—beneficial when the user is correct—but also intensifies regressive sycophancy when the user is wrong. For prompt design, this means that evidence-rich prompting should be used selectively: in contexts where the truth of the premise is already established, it can strengthen correct alignment; however, in ambiguous or high-stakes situations, models should be encouraged to independently verify evidence rather than simply align with it.
  4. 4.
    Framework Generalizability: Our progressive/regressive categorization and rebuttal chain evaluation framework provides a reusable methodology for measuring LLM reliability across domains. Because it focuses on the direction of model alignment and the strength of its corrective responses, this framework can be adapted for other high-stakes settings, such as law, finance, or engineering, where factual correctness must be maintained in the face of persuasive but incorrect user inputs
Limitations and Future Directions
The reliance on synthetic rebuttals may not fully capture real-world interaction diversity. Incorporating user-generated rebuttals could enhance generalizability. Additionally, our analysis focuses on three models; expanding this scope would provide broader insights. Finally, beta distribution modeling for LLM-as-a-Judge assumes consistent human evaluation, which warrants further investigation.
Future work should explore mitigating regressive sycophancy through hybrid reasoning architectures and longitudinal studies on retraining effects. Balancing alignment and truthfulness remains critical for deploying LLMs in high-stakes environments.
Conclusion
This study presents a comprehensive framework for assessing sycophantic behavior in LLMs, highlighting its dual nature and identifying factors influencing model behavior. These findings lay the groundwork for developing reliable AI systems for high-stakes applications, where accuracy must take precedence over user alignment.